# MathMathgg

### v1

@@ Appendix B Perturbative Expression of the RGTME
We write Eq. (29) more explicitly as,
@@
to describe the evolution of trial states under the non-perturbative Hamiltonian $H_{0}$. We also introduce
@@
as the first-order correction to Eq. (44). With Eq. (44) and (45), it easy to derive,
@@ Appendix C More about the Monte Carlo Method
In this work, we need to calculate matrix elements of the following form,
@@
as is shown in Eq. (46) and (47). Applying auxiliary field transformation, we have,
@@
Here $s$ is the auxiliary field located at all the lattice points and $S[s]$ is a Gaussian-type function. The interaction between the nucleons and the auxiliary field is included in $Z_{a,b}^{O}[s]$. After the transformation, both the nominators and the denominators in Eq. (46),(47) become integrals over $s$. We evaluate them through generating a set of auxiliary fields which distribute according to a given probability distribution function $P[s]$. A proper choice of $P[s]$ can help improve the sampling efficiency and reduce statistical noises. As both the initial and the final state nucleus contribute to the $\beta$-decay process, we take the following form of $P[s]$:

### v2

@@ Appendix B Perturbative Expression of the RGTME
We write Eq. (27) more explicitly as
@@
to describe the evolution of trial states under the nonperturbative Hamiltonian $H_{0}$. We also introduce
@@
as the first-order correction to Eq. (43). With Eq. (43) and (44), it easy to derive
@@ Appendix C More about the Monte Carlo Method
In this work, we need to calculate matrix elements of the following form
@@
as is shown in Eqs. (45) and (46). Applying auxiliary field transformation, we have,
@@
Here $s$ is the auxiliary field located at all the lattice points and $S[s]$ is a Gaussian-type function. The interaction between the nucleons and the auxiliary field is included in $Z_{a,b}^{O}[s]$. After the transformation, both the nominators and the denominators in Eqs. (45) and (46) become integrals over $s$. We evaluate them through generating a set of auxiliary fields which distribute according to a given probability distribution function $P[s]$. A proper choice of $P[s]$ can help improve the sampling efficiency and reduce statistical noises. As both the initial and final states of the nucleus nucleus contribute to the $\beta$-decay process, we take the following form of $P[s]$: